ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 47, 30 Hornor, William E., University of Southern Mississippi, USA Properties of isometry-inducing maps of the unit disc. We examine valence properties of self maps φ of the unit disc which induce isometries on non-Hilbertian Bergman or Hardy spaces. All isometries of these spaces are given by weighted composition operators T f = ω·f ∘φ, and we isolate properties of the inducing map φ. We examine the relationship between φ and the multiplier ω in the Bergman space setting. This talk is based on joint work with James E. Jamison.